The critical temperature Tc for Bose Einstein condensation varies with the number of density of particle as
1)n^(1/3)
2)n^(2/3)
3)n
4)n^(4/3)
1
Expert's answer
2019-07-04T11:12:50-0400

for critical temperature of Bose Einstein condensation we have equation

"T_c=(\frac{n}{\zeta(3/2)})^{\frac{2}{3}}\cdot \frac{2\pi h^2}{mk_b}"

"T_c \sim n^{\frac{2}{3}}"